OpenStudy (anonymous):

Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation. Part B: How many liters of milk are present in the bottle? Show your work.

OpenStudy (anonymous):

a: so if 20% is milk, and the 'rest' is water, then let y = total amount of water in liter and x = milk y = 1-.2*x the '1' is there because the bottle makes up 100% and since milk makes up 20% we know that 80% of the bottle must be water. now if we wanted an equation that would tell us the amount of milk in the bottle, as opposed to how much water is in the litter, then we would write it as x = 1-.8y

OpenStudy (ranga):

The problem states that only one variable can be used. Let L be the total number of liters of water and milk combined. 20% of the total is Milk. Therefore, Milk = L x 20/100 = 0.2 L The rest of it is Water. So Water = L - 0.2L = 0.8L The problem states the bottle has 1 liter of water. So 0.8L = 1 Answer to a) 0.8L = 1 is the equation where L is the total number of liters of water and milk combined. Milk = 0.2L and Water = 0.8L b) Solve for L in 0.8L = 1 L = 1/0.8 = 1.25 Milk = 0.2L = (0.2)(1.25) = 0.25 liters

OpenStudy (anonymous):

Thanks so much that explains it perfectly @ranga

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

Ranga please stop helping people cheat. you just gave trytohelper the answer. How is she learning?

OpenStudy (anonymous):

Obviously, she is learning how to solve this equation by asking for help and she got it by a person answering her question.

OpenStudy (anonymous):

@ducky2850